The amount of an ordinary $7,500.00 annuity for 3 years at 12% compounded quarterly is:
$23,182.50
$25,305.00
$180,997.50
$106,440.00
none of the above

Answer by rfer(16322)

A=7500(1+0.12/4)^4*3
A=7500(1.03)^12
A=7500*1.42576
A=$10,693.20
None of the above